# On pro-cdh descent on derived schemes

### v1

@@ Abstract.
We prove a ‘pro-cdh descent’ result for suitably connective localizing invariants and the cotangent complex on arbitrary qcqs derived schemes. As an application, we deduce a generalised Weibel vanishing for negative $K$-groups of non-Noetherian schemes.
@@ 1. Introduction
is weakly cartesian and hence induces a long exact sequence of pro-abelian groups. Note here that the pro-systems only depend on the underlying topological space of $Z$ (and $E$) as a closed subset of $X$ (and $\operatorname{Bl}_{Z}(X)$, respectively) but not on the particular subscheme structure. In fact, more generally any abstract blowup square of Noetherian schemes gives rise to a weakly cartesian square of $K$-theory pro-spectra. In this form, this result, often coined “pro-cdh descent”, [^1] We remark that it would be more appropriate to call this “pro-cdh excision” since it is not precisely a descent statement for some topology. However, together with Nisnevich excision, it does imply actual (Čech) descent for the procdh topology of [KS24], see Theorem 6.1 there. For this reason, and to be in line with the existing literature we stick to the term “pro-cdh descent.” is first proven in [KST18b], but there are important precursors [KS02, Cor06, Kri10, GH06, GH11, Mor16a, Mor18]. A variant for Noetherian ANS stacks has been considered in [BKRS22]. These pro-cdh descent results play a central role in the resolution of Weibel’s $K$-dimension conjecture [KST18b] (and [BKRS22] for stacks) and the development of a continuous $K$-theory of rigid spaces [Mor16b, KST18a, KST23].
Given the rising interest in non-Noetherian schemes, which appear naturally, for example, when working over a perfectoid base ring such as $\mathcal{O}_{\mathbb{C}_{p}}$, it is an obvious question, whether, or in which form, pro-cdh descent also holds in this general setting. The following example, a variant of which was constructed by Dahlhausen and the third author [DT22] precisely for that purpose and which was independently studied by the first two authors [Kel24, Footnote 2] with respect to the procdh topology, shows that pro-cdh descent as formulated above does not hold for general abstract blowup squares of non-Noetherian schemes.
@@
A posteriori, this is not surprising: A key point in [KST18b] was to prove Thomason’s blowup formula for derived blowups in quasi-smooth centers together with the fact that for every closed subscheme of an affine Noetherian scheme there exists a quasi-smooth derived closed subscheme with the same underlying classical scheme. The condition that all (derived) rings and schemes appearing be Noetherian then allowed to pass from derived schemes back to classical schemes. For instance, for a Noetherian commutative ring $A$ and elements $f_{1},\dots,f_{r}$ in $A$ generating the ideal $I\subseteq A$, the pro-system of (Koszul type) derived quotients $\{A/\!\!/f_{1}^{n},\dots,f_{r}^{n}\}_{n}$ is equivalent to the pro-system of discrete rings $\{A/I^{n}\}_{n}$. This suggested that some variant of the above pro-cdh descent statement, where one takes “derived” infinitesimal thickenings everywhere, could still be true. This is precisely what we achieve in this paper. Our main $K$-theoretic result is the following.
@@
In [KST18b], pro-cdh descent of $K$-theory was used to derive Weibel’s conjecture on the vanishing of negative $K$-groups of a Noetherian scheme $X$ below $-\dim(X)$. It is an obvious question what we can say about negative $K$-groups if we merely assume $X$ to be qcqs; see e.g. Morrow’s Oberwolfach talk [Mor23]. In this generality, one can at least prove a vanishing result for Weibel’s homotopy $K$-theory $KH(X)$. Let $\operatorname{vdim}(X)$ denote the valuative dimension of a scheme $X$ introduced by Jaffard [Jaf60] (see [EHIK20] or [KS24, §7.1] for accounts). It coincides with the Krull dimension if $X$ is Noetherian. Furthermore, write $L_{\mathrm{cdh}}$ for the cdh sheafification functor on presheaves of spectra and $K_{\geq 0}$ for the presheaf of connective algebraic $K$-theory.
@@
Though we are mainly interested in (1), we include (2) for the sake of completeness. For schemes essentially of finite type over a field of characteristic 0, this was first proven in [Hae04]. For general Noetherian $X$, (1) was first proven in [KS17] and (2) in [KST18b]. The fact that $KH$ is a cdh sheaf is [Cis13]. In the general case, the theorem follows easily from recent results on the cdh-topology [EHIK20] and the $K$-theory of valuation rings [KM21, KST21]. In fact, the proof given under Noetherian assumptions in [KM21] still works, and for the reader’s convenience we reproduce this proof at the end of the paper. Alternatively, the proof of [KS17] respectively [KST18b] works mutatis mutandis, using the fact that a blowup does not increase the valuative dimension (this is not necessarily true for the Krull dimension).
The previous theorem suggests that for any qcqs scheme $X$, the negative $K$-groups vanish below $-\operatorname{vdim}(X)$. In this direction, we prove the following.
Theorem B (Theorem 6.1).
Let $X$ be a qcqs (derived) scheme. Assume that the underlying topological space of the spectrum of every local ring of every irreducible component of $X$ is Noetherian. Then the following hold.
  1. (1)
    $K_{-i}(X)=0$ for all $i>\operatorname{vdim}(X)$.
  2. (2)
Remarks 1.4.
  1. (1)
    The assumption is for instance satisfied for any qcqs derived scheme of Krull dimension $\leq 1$. It is also stable under passing to schemes essentially of finite type over $X$. In particular, the theorem holds for all schemes essentially of finite type over a valuation ring.
  2. (2)
    For general qcqs $\mathbb{Q}$-schemes, (1) was announced by Elmanto and Morrow [Mor23, Thm. 4]. Moreover, they prove that the canonical map $K_{-i}(X)\to KH_{-i}(X)$ is an isomorphism for $i\geq\dim(X)$. Our result implies this only for $i\geq\operatorname{vdim}(X)$. Conversely, their isomorphism does not imply our assertion (2).
  3. (3)
    We do not know if the valuative dimension is really needed in the vanishing bound. I.e., we don’t know an example of a scheme $X$ with $\operatorname{vdim}(X)>\dim(X)$ and $K_{-\operatorname{vdim}(X)}(X)\not=0$.
@@
The proof of Theorems A and B follows the outline of [KST18b], but requires substantially more input from derived geometry. The proof of Theorem A is reduced to the two special cases of derived blowups and finite morphisms, respectively. This reduction is achieved by a structural result about modifications of derived schemes, Theorem 3.2, which we view as our main contribution here. In its proof, we need several preliminary results from derived algebraic geometry which we discuss in Section 2.
The case of Theorem A for derived blowups could essentially be proved as in [KST18b]. We here present a stronger result with a simplified proof due to Antieau ([Ant18]; we reproduce his proof in Proposition 4.1). For the case of finite morphisms in Theorem A, compared to [KST18b] we give a simple proof of a more general result (Proposition 4.2). This follows quite directly from [LT19].
The proof of our generalised Weibel vanishing statement is contained in Section 6. As we are working with non-Noetherian schemes, this needs some additional input.
@@
In forthcoming work, the first two authors introduce a “procdh topology” on qcqs derived schemes. It will follow from Theorem A (or 4.5) that every localizing invariant which is $k$-connective for some integer $k$ satisfies descent for that topology.
@@ Acknowledgement
We thank Christian Dahlhausen for discussions about pro-cdh descent in the non-Noetherian setting which led to the variant of Example 1.1 and which was one of the starting points of this project. We are grateful to Ben Antieau for allowing us to include his pro-excision result for derived blowups, which is stronger than the one obtained in [KST18b]. We thank Matthew Morrow for bringing our attention to the cotangent complex and further comments and Mauro Porta for related discussions.
@@ 2.2. Quasi-coherent modules
If $Z\subseteq|X|$ is a closed subset, we denote by $\operatorname{QCoh}(X\,\mathrm{on}\,Z)$ the full subcategory of $\operatorname{QCoh}(X)$ spanned by those quasi-coherent sheaves which are supported on $Z$. If $X$ is quasi-compact and quasi-separated (qcqs, for short) and the open complement of $Z$ is compact, then $\operatorname{QCoh}(X\,\mathrm{on}\,Z)$ is compactly generated and its compact objects coincide with the perfect ones, [Lur18, Prop. 9.6.1.1], [CMNN, Prop. A.9]. Here a quasi-coherent sheaf on $X$ is called perfect if and only its restriction to each affine open subscheme $U=\operatorname{Spec}(A)\subseteq X$ belongs to the smallest thick stable subcategory of $\operatorname{QCoh}(U)=\operatorname{Mod}(A)$ containing $A$. We write $\operatorname{Perf}(X)$ and $\operatorname{Perf}(X\,\mathrm{on}\,Z)$ for the corresponding subcategories. The arguments used to prove this also apply to the case of connective sheaves:
@@
Let $X$ be a qcqs derived scheme, and let $Z\subseteq|X|$ be a closed subset with compact open complement. Then $\operatorname{QCoh}(X\,\mathrm{on}\,Z)^{\mathrm{cn}}$ is compactly generated and the inclusion $\operatorname{QCoh}(X\,\mathrm{on}\,Z)^{\mathrm{cn}}\hookrightarrow\operatorname{QCoh}(X)^{\mathrm{cn}}$ preserves compact objects. An object of $\operatorname{QCoh}(X\,\mathrm{on}\,Z)^{\mathrm{cn}}$ is compact if and only if it is perfect (as an object of $\operatorname{QCoh}(X)$).
@@
The same argument as in the proof of [CMNN, Prop. A.9] proves the first two claims: In case $X$ is affine, these follow from [Lur18, Prop. 7.1.1.12(e)] and the fact that $\operatorname{QCoh}(X)^{\mathrm{cn}}$ is compactly generated by [Lur18, Prop. 9.6.1.2]. The reduction of the global case to the local case is done by [Lur18, Ex. 10.3.0.2 (4), Prop. 10.3.0.3, Th. 10.3.2.1 (b)]. It then follows that an object of $\operatorname{QCoh}(X\,\mathrm{on}\,Z)^{\mathrm{cn}}$ is compact if and only if its image in $\operatorname{QCoh}(X)^{\mathrm{cn}}$ is compact. The compact objects in the latter category coincide with the perfect, connective $\mathcal{O}_{X}$-modules by [Lur18, Prop. 9.6.1.2.] again. ∎
@@ 2.4. Finiteness conditions
These finiteness conditions are stable under base change: If $A\to B$ is locally or almost of finite presentation and $A\to A^{\prime}$ is an arbitrary map, then also $A^{\prime}\to B\otimes_{A}A^{\prime}$ is locally or almost of finite presentation, respectively [Lur17, Rem. 7.2.4.28]. They are also stable under composition [Lur17, Rem. 7.2.4.29, Cor. 7.4.3.19].
@@
Let $X$ be a qcqs derived scheme, and let $U\subseteq X$ be a qc open subset with complement $Z$. Let $i\colon Y\to X$ be closed immersion which is an isomorphism over $U$ and such that the underlying map of classical schemes $\prescript{\mathrm{cl}}{}{Y}\to\prescript{\mathrm{cl}}{}{X}$ is finitely presented. Then there exists a factorization $Y\to Y^{\prime}\to X$ of $i$ such that
@@ 2.6. Ample line bundles
[…]
@@ 2.7. Quasi-smooth closed immersions and derived blowups
Derived blowups were first introduced in [KST18b] for affine schemes in order to prove pro-cdh descent of algebraic $K$-theory on Noetherian schemes. They were then systematically studied and developed much further by Khan, Rydh, and Hekking [KR18, Hek21].
Let $X$ be a derived scheme, and let $Z\hookrightarrow X$ be a quasi-smooth closed immersion [KR18, 2.3.6], i.e., Zariski locally on $X$, $Z\hookrightarrow X$ is the derived pullback of the map $\{0\}\hookrightarrow\mathbb{A}^{r}_{\mathbb{Z}}$ for some $r$ and some morphism $X\to\mathbb{A}^{r}_{\mathbb{Z}}$. Equivalently, Zariski locally on $X$, $Z\hookrightarrow X$ is given by $\operatorname{Spec}(A/\!\!/f_{1},\dots,f_{r})\hookrightarrow\operatorname{Spec}(A)$ for suitable elements $f_{i}\in\pi_{0}(A)$. The number $r$ is called the virtual codimension of the closed immersion. Then one can form the derived blowup $p\colon\operatorname{dBl}_{Z}(X)\to X$ of $X$ in $Z$ (or with center $Z$) which is characterized by a universal property: it classifies virtual Cartier divisors on derived $X$-schemes; see [KR18, §4.1] for details. The construction of the derived blowup commutes with arbitrary base change. Locally, if $Z$ is the derived pullback of $\{0\}\hookrightarrow\mathbb{A}^{r}_{\mathbb{Z}}$ along a map $X\to\mathbb{A}^{r}_{\mathbb{Z}}$, then $\operatorname{dBl}_{Z}(X)$ is the derived pullback of the classical blowup $\operatorname{Bl}_{\{0\}}(\mathbb{A}^{r}_{\mathbb{Z}})\to\mathbb{A}^{r}_{\mathbb{Z}}$. By the discussion of finiteness conditions in 2.4, as $\mathbb{A}^{r}_{\mathbb{Z}}$ is Noetherian, this in particular implies that the morphism $p\colon\operatorname{dBl}_{Z}(X)\to X$ is lafp. Clearly, $p$ is an isomorphism outside $Z$ and proper.
@@
The derived blowup $\operatorname{dBl}_{Z}(X)$ carries a canonical line bundle $\mathcal{O}(1)$ (the dual of the ideal sheaf defining the universal virtual Cartier divisor) and this line bundle is $p$-ample. Indeed, by Lemma 2.4 we may work affine-locally on $X$ and hence assume that $p\colon\operatorname{dBl}_{Z}(X)\to X$ is the pullback of the classical blowup $\operatorname{Bl}_{\{0\}}(\mathbb{A}^{n}_{\mathbb{Z}})\to\mathbb{A}^{n}_{\mathbb{Z}}$, and the line bundle $\mathcal{O}(1)$ is the pullback of the classical canonical line bundle on $\operatorname{Bl}_{\{0\}}(\mathbb{A}^{n})$ which is ample. As being relatively ample is stable under base change (the proof in [Ann22, Prop. 2.12] works in general), it follows that $\mathcal{O}(1)$ is $p$-ample.

### v2

@@ Abstract.
We prove a ‘pro-cdh descent’ result for suitably connective localizing invariants and the cotangent complex on arbitrary qcqs derived schemes. As an application, we deduce that the $K$-groups of arbitrary qcqs schemes vanish below the negative of their valuative dimension.
@@ 1. Introduction
is weakly cartesian and hence induces a long exact sequence of pro-abelian groups. Note here that the pro-systems only depend on the underlying topological space of $Z$ (and $E$) as a closed subset of $X$ (and $\operatorname{Bl}_{Z}(X)$, respectively) but not on the particular subscheme structure. In fact, more generally any abstract blowup square of Noetherian schemes gives rise to a weakly cartesian square of $K$-theory pro-spectra. In this form, this result, often coined “pro-cdh descent”, [^1] We remark that it would be more appropriate to call this “pro-cdh excision” since it is not precisely a descent statement for some topology. However, together with Nisnevich excision, it does imply actual (Čech) descent for the pro-cdh topology of [KS24], see Theorem 6.1 there. For this reason, and to be in line with the existing literature we stick to the term “pro-cdh descent.” is first proven in [KST18], but there are important precursors [KS02, Cor06, Kri10, GH06, GH11, Mor16b, Mor18]. A variant for Noetherian ANS stacks has been considered in [BKRS22]. These pro-cdh descent results play a central role in the resolution of Weibel’s $K$-dimension conjecture [KST18] (and [BKRS22] for stacks) and the development of a continuous $K$-theory of rigid spaces [Mor16a, KST19, KST23].
Given the rising interest in non-Noetherian schemes, which appear naturally, for example, when working over a perfectoid base ring such as $\mathcal{O}_{\mathbb{C}_{p}}$, it is an obvious question, whether, or in which form, pro-cdh descent also holds in this general setting. The following example, a variant of which was constructed by Dahlhausen and the third author [DT22] precisely for that purpose and which was independently studied by the first two authors [Kel24, Footnote 2] with respect to the pro-cdh topology, shows that pro-cdh descent as formulated above does not hold for general abstract blowup squares of non-Noetherian schemes.
@@
A posteriori, this is not surprising: A key point in [KST18] was to prove Thomason’s blowup formula for derived blowups in quasi-smooth centers together with the fact that for every closed subscheme of an affine Noetherian scheme there exists a quasi-smooth derived closed subscheme with the same underlying classical scheme. The condition that all (derived) rings and schemes appearing be Noetherian then allowed to pass from derived schemes back to classical schemes. For instance, for a Noetherian commutative ring $A$ and elements $f_{1},\dots,f_{r}$ in $A$ generating the ideal $I\subseteq A$, the pro-system of (Koszul type) derived quotients $\{A/\!\!/f_{1}^{n},\dots,f_{r}^{n}\}_{n}$ is equivalent to the pro-system of discrete rings $\{A/I^{n}\}_{n}$. This suggested that some variant of the above pro-cdh descent statement, where one takes “derived” infinitesimal thickenings everywhere, could still be true. This is precisely what we achieve in this paper. Our main $K$-theoretic result is the following.
@@
In [KST18], pro-cdh descent of $K$-theory was used to derive Weibel’s conjecture on the vanishing of negative $K$-groups of a Noetherian scheme $X$ below $-\dim(X)$. It is an obvious question what we can say about negative $K$-groups if we merely assume $X$ to be qcqs; see e.g. Morrow’s Oberwolfach talk [Mor23]. As the Krull dimension doesn’t have good properties for general non-Noetherian schemes, one might expect that instead the valuative dimension introduced by Jaffard [Jaf60] (see [EHIK20] or [KS24, §7.1] for accounts) gives the correct bounds. It coincides with the Krull dimension if $X$ is Noetherian. Indeed, we prove the following.
Theorem B (Theorem 6.5).
Let $X$ be a qcqs spectral scheme. Then the following hold.
  1. (1)
    $K_{-i}(X)=0$ for all $i>\operatorname{vdim}(X)$.
  2. (2)
Here by a spectral scheme we mean more precisely one with connective structure sheaf, and in (2), $\mathbb{A}^{r}$ denotes either the flat or the smooth affine space.
Remark 1.3.
We don’t know an example of a scheme $X$ with $\operatorname{vdim}(X)>\dim(X)$ and $K_{-\operatorname{vdim}(X)}(X)\not=0$.
This theorem immediately implies the same vanishing for Weibel’s homotopy $K$-theory $KH(X)$. In other words, we get a new proof of assertion (1) in the following theorem, which we include here for the sake of completeness. As we discuss below, this theorem is well known to the experts. We write $L_{\mathrm{cdh}}$ for the cdh sheafification functor on presheaves of spectra and $K_{\geq 0}$ for the presheaf of connective algebraic $K$-theory.
@@
For schemes essentially of finite type over a field of characteristic 0, this was first proven in [Hae04]. For general Noetherian $X$, (1) was first proven in [KS17] and (2) in [KST18]. The fact that $KH$ is a cdh sheaf is [Cis13]. In the general case, the theorem follows easily from recent results on the cdh-topology [EHIK20] and the $K$-theory of valuation rings [KM21, KST21]. In fact, the proof given under Noetherian assumptions in [KM21] still works, and for the reader’s convenience we reproduce this proof at the end of the paper. Alternatively, the proof of [KS17] respectively [KST18] works mutatis mutandis, using the fact that a blowup does not increase the valuative dimension (this is not necessarily true for the Krull dimension).
Remark 1.5.
For arbitrary qcqs schemes $X$, Theorem 1.4 together with the cartesian square
${K}$${KH}$${\mathrm{TC}}$${L_{cdh}\mathrm{TC},}$
$(-1)$-connectivity of topological cyclic homology on affines, and [EHIK20] imply the slightly weaker vanishing $K_{-i}(X)=0$ for $i>\operatorname{vdim}(X)+2$. Elmanto and Morrow (over fields) and Bouis (in general) prove refinements of this for the motivic filtration [EM23, Proof of Thm. 4.12], [Bou24, Prop. 5.5.4].
@@
The proof of Theorem A follows the outline of [KST18], but requires substantially more input from derived geometry. It is reduced to the two special cases of derived blowups and finite morphisms, respectively. This reduction is achieved by a structural result about modifications of derived schemes, Theorem 3.2, which we view as our main contribution here. In its proof, we need several preliminary results from derived algebraic geometry which we discuss in Section 2.
The case of Theorem A for derived blowups could essentially be proved as in [KST18]. We here present a stronger result with a simplified proof due to Antieau ([Ant18]; we reproduce his proof in Proposition 4.1). For the case of finite morphisms in Theorem A, compared to [KST18] we give a simple proof of a more general result (Proposition 4.2).
To prove Theorem B in Section 6, we roughly follow Kerz’s approach [Ker18] combined with our pro-cdh-descent for finite morphisms. Another new input is a generalization of results of Swanson and Huneke [HS06] on reductions of ideals to the non-Noetherian setting in Subsection 6.2.
@@
In forthcoming work, the first two authors introduce a “pro-cdh topology” on qcqs derived schemes. It will follow from Theorem A (or 4.6) that every localizing invariant which is $k$-connective for some integer $k$ satisfies descent for that topology.
@@ Acknowledgement
We thank Christian Dahlhausen for discussions about pro-cdh descent in the non-Noetherian setting which led to the variant of Example 1.1 and which was one of the starting points of this project. We are grateful to Ben Antieau for allowing us to include his pro-excision result for derived blowups, which is stronger than the one obtained in [KST18]. We thank Matthew Morrow for bringing our attention to the cotangent complex and further comments and Mauro Porta for related discussions. Finally we thank David Rydh for enlightening comments on our Sections 2 and 3. In particular, he proposed to formulate the more precise version of Theorem 3.2 including different finiteness hypotheses.
@@ 2.2. Quasi-coherent modules
If $Z\subseteq|X|$ is a closed subset, we denote by $\operatorname{QCoh}(X\,\mathrm{on}\,Z)$ the full subcategory of $\operatorname{QCoh}(X)$ spanned by those quasi-coherent sheaves which are supported on $Z$. If $X$ is quasi-compact and quasi-separated (qcqs, for short) and the open complement of $Z$ is quasi-compact, then $\operatorname{QCoh}(X\,\mathrm{on}\,Z)$ is compactly generated and its compact objects coincide with the perfect ones, [Lur18, Prop. 9.6.1.1], [CMNN20, Prop. A.9]. Here a quasi-coherent sheaf on $X$ is called perfect if and only its restriction to each affine open subscheme $U=\operatorname{Spec}(A)\subseteq X$ belongs to the smallest thick stable subcategory of $\operatorname{QCoh}(U)=\operatorname{Mod}(A)$ containing $A$. We write $\operatorname{Perf}(X)$ and $\operatorname{Perf}(X\,\mathrm{on}\,Z)$ for the corresponding subcategories. The arguments used to prove this also apply to the case of connective sheaves:
@@
Let $X$ be a qcqs derived scheme, and let $Z\subseteq|X|$ be a closed subset with quasi-compact open complement. Then $\operatorname{QCoh}(X\,\mathrm{on}\,Z)^{\mathrm{cn}}$ is compactly generated and the inclusion $\operatorname{QCoh}(X\,\mathrm{on}\,Z)^{\mathrm{cn}}\hookrightarrow\operatorname{QCoh}(X)^{\mathrm{cn}}$ preserves compact objects. An object of $\operatorname{QCoh}(X\,\mathrm{on}\,Z)^{\mathrm{cn}}$ is compact if and only if it is perfect (as an object of $\operatorname{QCoh}(X)$).
@@
The same argument as in the proof of [CMNN20, Prop. A.9] proves the first two claims: In case $X$ is affine, these follow from [Lur18, Prop. 7.1.1.12(e)] and the fact that $\operatorname{QCoh}(X)^{\mathrm{cn}}$ is compactly generated by [Lur18, Prop. 9.6.1.2]. The reduction of the global case to the local case is done by [Lur18, Ex. 10.3.0.2 (4), Prop. 10.3.0.3, Th. 10.3.2.1 (b)]. It then follows that an object of $\operatorname{QCoh}(X\,\mathrm{on}\,Z)^{\mathrm{cn}}$ is compact if and only if its image in $\operatorname{QCoh}(X)^{\mathrm{cn}}$ is compact. The compact objects in the latter category coincide with the perfect, connective $\mathcal{O}_{X}$-modules by [Lur18, Prop. 9.6.1.2.] again. ∎
@@ 2.4. Finiteness conditions
For completeness, we also mention that for an integer $n\geq 0$, there is a notion of finite generation to order n, see [Lur18, Def. 4.1.1.1], and a morphism is almost finitely presented if and only if it is of finite generation to order $n$ for all $n$.
These finiteness conditions are stable under base change: If $A\to B$ is locally or almost of finite presentation or of finite generation to order $n$ and $A\to A^{\prime}$ is an arbitrary map, then also $A^{\prime}\to B\otimes_{A}A^{\prime}$ is locally or almost of finite presentation or of finite generation to order $n$, respectively [Lur17, Rem. 7.2.4.28], [Lur18, Prop. 4.1.3.2]. They are also stable under composition [Lur17, Rem. 7.2.4.29, Cor. 7.4.3.19], [Lur18, Prop. 4.1.3.1].
@@
Let $X$ be a qcqs derived scheme, and let $U\subseteq X$ be a qc open subset with complement $Z$. Let $i\colon Y\to X$ be a closed immersion which is an isomorphism over $U$ and such that the underlying map of classical schemes $\prescript{\mathrm{cl}}{}{Y}\to\prescript{\mathrm{cl}}{}{X}$ is finitely presented. Then there exists a factorization $Y\to Y^{\prime}\to X$ of $i$ such that
@@ 2.6. Ample line bundles
$\operatorname*{colim}_{n}\Gamma(X,\mathcal{F}\otimes\mathcal{L}^{\otimes n})\to\operatorname*{colim}_{n}\Gamma(X_{s},\mathcal{F}\otimes\mathcal{L}^{\otimes n})\simeq\Gamma(X_{s},\mathcal{F}).$
@@ 2.7. Quasi-smooth closed immersions and derived blowups
Derived blowups were first introduced in [KST18] for affine schemes in order to prove pro-cdh descent of algebraic $K$-theory on Noetherian schemes. They were then systematically studied and developed much further by Khan, Rydh, and Hekking [KR18, Hek21].
Let $X$ be a derived scheme, and let $Z\hookrightarrow X$ be a quasi-smooth closed immersion [KR18, 2.3.6], i.e., Zariski locally on $X$, $Z\hookrightarrow X$ is the derived pullback of the map $\{0\}\hookrightarrow\mathbb{A}^{r}_{\mathbb{Z}}$ for some $r$ and some morphism $X\to\mathbb{A}^{r}_{\mathbb{Z}}$. Equivalently, Zariski locally on $X$, $Z\hookrightarrow X$ is given by $\operatorname{Spec}(A/\!\!/f_{1},\dots,f_{r})\hookrightarrow\operatorname{Spec}(A)$ for suitable elements $f_{i}\in\pi_{0}(A)$. The number $r$ is called the virtual codimension of the closed immersion. Then one can form the derived blowup $p\colon\operatorname{dBl}_{Z}(X)\to X$ of $X$ in $Z$ (or with center $Z$) which is characterized by a universal property: it classifies virtual Cartier divisors on derived $X$-schemes; see [KR18, §4.1] for details. The construction of the derived blowup commutes with arbitrary base change. Locally, if $Z$ is the derived pullback of $\{0\}\hookrightarrow\mathbb{A}^{r}_{\mathbb{Z}}$ along a map $X\to\mathbb{A}^{r}_{\mathbb{Z}}$, then $\operatorname{dBl}_{Z}(X)$ is the derived pullback of the classical blowup $\operatorname{Bl}_{\{0\}}(\mathbb{A}^{r}_{\mathbb{Z}})\to\mathbb{A}^{r}_{\mathbb{Z}}$. By [KR18, Thm. 4.1.5(v)], the morphism $p\colon\operatorname{dBl}_{Z}(X)\to X$ is quasi-smooth, i.e. Zariski locally factors as a quasi-smooth closed immersion followed by a smooth morphism, and in particular is locally of finite presentation. Clearly, $p$ is an isomorphism outside $Z$ and proper.
@@
The derived blowup $\operatorname{dBl}_{Z}(X)$ carries a canonical line bundle $\mathcal{O}(1)$ (the ideal sheaf defining the universal virtual Cartier divisor) and this line bundle is $p$-ample. Indeed, by Lemma 2.4 we may work affine-locally on $X$ and hence assume that $p\colon\operatorname{dBl}_{Z}(X)\to X$ is the pullback of the classical blowup $\operatorname{Bl}_{\{0\}}(\mathbb{A}^{n}_{\mathbb{Z}})\to\mathbb{A}^{n}_{\mathbb{Z}}$, and the line bundle $\mathcal{O}(1)$ is the pullback of the classical canonical line bundle on $\operatorname{Bl}_{\{0\}}(\mathbb{A}^{n})$ which is ample. As being relatively ample is stable under base change (the proof in [Ann22, Prop. 2.12] works in general), it follows that $\mathcal{O}(1)$ is $p$-ample.